# A Public and Reproducible Assessment of the Topics API on Real Data

### v1

@@ Abstract
The Topics API for the web is Google’s privacy-enhancing alternative to replace third-party cookies. Results of prior work have led to an ongoing discussion between Google and research communities about the capability of Topics to trade off both utility and privacy. The central point of contention is largely around the realism of the datasets used in these analyses and their reproducibility; researchers using data collected on a small sample of users or generating synthetic datasets, while Google’s results are inferred from a private dataset. In this paper, we complement prior research by performing a reproducible assessment of the latest version of the Topics API on the largest and publicly available dataset of real browsing histories. First, we measure how unique and stable real users’ interests are over time. Then, we evaluate if Topics can be used to fingerprint the users from these real browsing traces by adapting methodologies from prior privacy studies. Finally, we call on web actors to perform and enable reproducible evaluations by releasing anonymized distributions. We find that , , and of the users in the dataset are uniquely re-identified across websites after only , , and observations of their topics by advertisers, respectively. This paper shows on real data that Topics does not provide the same privacy guarantees to all users, further highlighting the need for public and reproducible evaluations of the claims made by new web proposals.
@@ 1 Introduction
These results have led to an ongoing discussion between Google and research communities about the capability of the Topics API to deliver on both its utility and privacy objectives. The major point of contention between the analyses carried out by Google and researchers is the access asymmetry to real browsing data as well as the resulting reproducibility of the evaluations. Indeed, while researchers have either collected browsing data on a small sample of users [5] or synthetically generated large traces [6], Google performed their evaluations on a private dataset [7, 8] and only reported aggregate results [4], making it impossible to reproduce their evaluation.
@@
First, we measure on an anonymized dataset of over a month of real browsing histories how stable and unique users’ online behaviors and interests are over time. This is to compare with a stability assumption assumed in prior work [6]. Then, we adapt prior privacy analyses of the Topics API, but on the latest version of the proposal111Commit sha 78b13c3 as of February 19, 2024 as a new topics taxonomy, a new machine learning classifier, and other modifications of the topics computation have been released since. Finally, we simulate how the Topics API would work for these users and evaluate if advertisers can fingerprint and re-identify users though their topics.
We find that each week at least of the users in this dataset have a unique top topics profile. Measuring how stable these top profiles are over time, we obtain that at least of users have or more topics in common in their top topics profile from one week to the next one, while only less than of users have none. The combination of unique and stable profiles opens the possibility to use the Topics API to fingerprint users. We observe that in practice for the real users from the dataset, , , and of them are re-identified across websites after , , and observations of their topics by third parties, respectively. This paper highlights the importance of public and reproducible evaluations of any claim made by new web proposals and to identify the potential limitations of these techniques during their design rather than after their deployment.
@@ 2.1 Topics API for the Web
Description. The Topics API for the web aims at deprecating third-party cookies while still enabling interest-based advertising [1, 11, 2]. Google proposes that without third-party cookies, web browsers be the ones to classify users’ online behaviors into topics of interests and return the top visited topics to advertisers embedded on the websites being browsed. As of February 2024, the Topics API for the web is being rolled out to every Chrome users [12, 13], while other browsers such as Apple Safari [14], Mozilla Firefox [15, 4], or Brave [16] have declined to implement Topics in their product citing several privacy concerns they have.
@@ 3.1 New Topics API Version
Modifications to Domains Classification. First, a new taxonomy (version 2) of topics has been released: topics from the initial taxonomy have been removed and new ones added. The model classifier (version 4) has been updated accordingly, and the static mapping that consists of a list of websites to topics manually annotated by Google has grown from about k to now k domains.
@@
Re-implementation and Topics Distribution. For this paper, we re-implement the latest version of the Topics API (commit sha 78b13c3) shipped in Google Chrome locally to be able to classify several thousands domains at once. This lets us obtain the prior distributions of topics observations on the most visited websites; Figure 2 shows such distributions on the following three lists of domains: the new static mapping used by Google composed of about 50k domains (2(a)), and the top 1M most visited domains as reported by CrUx [25] (2(b)) and Tranco [26] (2(c)), respectively. While the Topics API has been updated since prior analyses, results about the asymmetry of the observations of topics from the new taxonomy on the most visited websites still hold [6]; , , and topics never appear at all from the taxonomy for the static mapping, CrUX, and Tranco, respectively, others only a few times, and very few on many domains from each list. Such prior distribution can be used by third party to distinguish if an observed topic is a noisy one (picked randomly from the taxonomy by the API in of the cases) if it never appears or only seldomly on the top M domains in practice (see Section 4.4 for more).
@@ 4.3 Stability and Uniqueness of Topics Profiles
From week 1 to 2 | 57 (4.7%) | 184 (15.2%) | 301 (24.9%) | 373 (30.9%) | 229 (19.0%) | 63 (5.2%) |
From week 2 to 3 | 67 (5.6%) | 193 (16.0%) | 315 (26.1%) | 353 (29.2%) | 220 (18.2%) | 59 (4.9%) |
From week 3 to 4 | 70 (5.8%) | 188 (15.6%) | 318 (26.3%) | 333 (27.6%) | 238 (19.7%) | 60 (5.0%) |
From week 4 to 5 | 70 (5.8%) | 233 (19.3%) | 329 (27.3%) | 317 (26.3%) | 215 (17.8%) | 43 (3.6%) |
@@
Unique topics | 218 | 215 | 220 | 223 | 226 |
Unique profiles | 1132 | 1132 | 1144 | 1145 | 1151 |
@@ 4.5 Re-identifying Users Across Sites
Finally, to respond to (Q3), we run the following re-identification experiment [6, 8]. We consider that each user visits two different websites on which two different advertisers that are colluding are embedded. For each week that we simulate, these advertisers are observing the topics of each user from their respective vantage points. Their objective is to correlate a user seen by one advertiser on the first website with the correct user seen by the other advertiser on the second website. For each user, this matching is done by returning the user(s) that have the lowest Hamming distance between the topics observed by the second advertiser and the set of topics observed by the first advertiser [8]. We leave to future work the exploration of other matching strategies. Also, this re-identification is performed every week with only the topics that have been observed until that week for each user. We present the results of this experiment in terms of level of -anonymity that is provided to users by the API; i.e.,, means that advertisers have been able to uniquely re-identify the user across their visits for that specific week, where is the number of users means that the user’s visits could not be uniquely correlated to each other, but with the visits of other users, and means that the re-identification failed and is no better than a random uniform guess. Figure 3 shows the cumulative distribution function of the proportion of users that are satisfied -anonymity (log scale) across the simulated weeks (recall that a topic observation for a user requires knowing their top interests for the last weeks, thus we start at week 3). Note that the experiment is also valid, and our results exactly the same, if we were to consider instead of two advertisers only one embedded on both websites and trying to correlate visits from both vantage points. We observe that , , and of the users are uniquely re-identified across their visits after only , , and observations, respectively. For the weeks simulated, we observe that about of the users are mapped to a group of size , which means that their cross-site visits can be correlated to each other with a higher probability than just a random guess. We conclude that an important part of the users from this real dataset are re-identified across websites through only the observations of their topics of interest in our experiment. Thus, the real users from our dataset can be fingerprinted through the Topics API. Moreover, as can be seen, the information leakage and so, privacy violation worsen over time as more users are uniquely re-identified.
@@ 5 Discussion
Future Work. Overall, rigorous and thorough analyses of these new web proposals (The Privacy Sandbox for the web now encompasses more than 30 different mechanisms [33]) are needed; several interesting questions remain unsolved, such as if these proposals can deliver on their guarantees to all users, if their specification and implementation are correct, if they can be manipulated or abused, if their results can be correlated with other fingerprinting signals, if developers and users do understand how these new APIs work, etc. Regarding the Topics API, evaluating it without having access to Google’s private dataset of browsing histories can be quite challenging; we refer readers to previous analyses of the Topics API to see how prior authors (partially) got around that obstacle [5, 6]. In this paper, we call on web actors to release anonymized or synthetic datasets to help with that point. In the meantime, we envision to perform a literature review and a taxonomy of web users measurement studies, pair these results with the few datasets of browsing histories publicly available, and generate synthetic distributions of users’ topics of interest.
@@ 6 Conclusion
This paper complements previous analyses of the Topics API for the web by confirming on real data that the proposal does not provide the same privacy guarantees to all users. Specifically, we have shown on this real dataset of users that the Topics API can be used to fingerprint of these users across websites after observations. We also highlighted the need for reproducible analyses of the claims made by new web proposals before their deployment. This is important to avoid replicating the same mistakes already made in the past with the same technologies being deprecated today; the introduction of e.g., web cookies later extensively abused for tracking purposes. As such, we call on web actors and industry to release representative and anonymized datasets to enable reproducible evaluations of the Topics API and other proposals from The Privacy Sandbox.
@@ Acknowledgments
We sincerely thank the reviewers for their constructive comments and suggestions on this paper.
@@ Appendix A Notations
Commit on Topics proposal | Commit sha of the Topics proposal studied in this paper | 78b13c3 (February 19, 2024) |
@@
CrUX version | Version of the CrUX top-list used in this paper | 202401 (January 2024) |
Tranco version | Version of the Tranco top-list used in this paper | 6JZJX (February 4, 2024) |

### v3

@@ Abstract
The Topics API for the web is Google’s privacy-enhancing alternative to replace third-party cookies. Results of prior work have led to an ongoing discussion between Google and research communities about the capability of Topics to trade off both utility and privacy. The central point of contention is largely around the realism of the datasets used in these analyses and their reproducibility; researchers using data collected on a small sample of users or generating synthetic datasets, while Google’s results are inferred from a private dataset. In this paper, we complement prior research by performing a reproducible assessment of the latest version of the Topics API on the largest and publicly available dataset of real browsing histories. First, we measure how unique and stable real users’ interests are over time. Then, we evaluate if Topics can be used to fingerprint the users from these real browsing traces by adapting methodologies from prior privacy studies. Finally, we call on web actors to perform and enable reproducible evaluations by releasing anonymized distributions. We find that for the real users in this dataset, the probability of being re-identified across websites is of , , and after , , and observations of their topics by advertisers, respectively. This paper shows on real data that Topics does not provide the same privacy guarantees to all users and that the information leakage worsens over time, further highlighting the need for public and reproducible evaluations of the claims made by new web proposals.
@@ 1 Introduction
These results have led to an ongoing discussion between Google and research communities about the capability of the Topics API to deliver on both its utility and privacy objectives. The major point of contention between the analyses carried out by Google and researchers is the access asymmetry to real browsing data as well as the resulting reproducibility of the evaluations. Indeed, while researchers have either collected browsing data on a small sample of users [5] or synthetically generated large traces [6], Google performed their evaluations on a private dataset [7, 8] and only reported aggregate results [4], making it impossible to reproduce Google’s evaluation.
@@
First, we measure on an anonymized dataset of over a month of real browsing histories how stable and unique users’ online behaviors and interests are over time. This is to compare with a stability assumption assumed in prior work [6]. Then, we adapt prior privacy analyses of the Topics API, but on the latest version of the proposal111Commit sha 2df537b as of July 8, 2024 as a new topics taxonomy, a new machine learning classifier, and other modifications of the topics computation have been released since. Finally, we simulate how the Topics API would work for these users and evaluate if advertisers can fingerprint and re-identify users through their topics.
We find that each week at least of the users in this dataset have a unique top topics profile. Measuring how stable these top profiles are over time, we obtain that at least of users have or more topics in common in their top topics profile from one week to the next one, while only less than of users have none. The combination of unique and stable profiles opens the possibility to use the Topics API to fingerprint users. We observe that in practice for the real users from the dataset, their probability of being re-identified across websites is of , , and after , , and observations of their topics by third parties, respectively. This paper highlights the importance of public and reproducible evaluations of any claim made by new web proposals and to identify the potential limitations of these techniques during their design rather than after their deployment.
@@ 2.1 Topics API for the Web
Description. The Topics API for the web aims at deprecating third-party cookies while still enabling interest-based advertising [1, 11, 2]. Google proposes that without third-party cookies, web browsers be the ones to classify users’ online behaviors into topics of interests and return the top visited topics to advertisers embedded on the websites being browsed. As of February 2024, the Topics API for the web is being rolled out to every Chrome users [12, 13], while other browsers such as Apple Safari [14], Mozilla Firefox [15, 4], or Brave [16] have declined to implement Topics in their product citing several privacy concerns with the proposal.
@@ 3.1 New Topics API Version
Modifications to Domains Classification. First, a new taxonomy (version 2) of topics has been released: topics from the initial taxonomy have been removed and new ones added. The model classifier (version 5) has been updated accordingly, and the static mapping that consists of a list of websites to topics manually annotated by Google has grown from about k to now k domains.
@@
Re-implementation and Topics Distribution. For this paper, we re-implement the latest version of the Topics API (commit sha 78b13c3) shipped in Google Chrome locally to be able to classify several thousands domains at once. This lets us obtain the prior distributions of topics observations on the most visited websites; Figure 2 shows such distributions on the following three lists of domains: the new static mapping used by Google composed of about 50k domains (2(a)), and the top 1M most visited domains as reported by CrUX [25] (2(b)) and Tranco [26] (2(c)), respectively. While the Topics API has been updated since prior analyses, results about the asymmetry of the observations of topics from the new taxonomy on the most visited websites still hold [6]; , , and topics never appear at all from the taxonomy for the static mapping, CrUX, and Tranco, respectively, others only a few times, and very few on many domains from each list. Such prior distribution can be used by third party to distinguish if an observed topic is a noisy one (picked randomly from the taxonomy by the API in of the cases) if it never appears or only seldomly on the top M domains in practice (see Section 4.4).
@@ 4.3 Stability and Uniqueness of Topics Profiles
From week 1 to 2 | 59 (4.9%) | 187 (15.5%) | 297 (24.6%) | 369 (30.6%) | 229 (19.0%) | 66 (5.5%) |
From week 2 to 3 | 70 (5.8%) | 189 (15.7%) | 318 (26.3%) | 345 (28.6%) | 226 (18.7%) | 59 (4.9%) |
From week 3 to 4 | 72 (6.0%) | 188 (15.6%) | 320 (26.5%) | 325 (26.9%) | 246 (20.4%) | 56 (4.6%) |
From week 4 to 5 | 70 (5.8%) | 240 (19.9%) | 324 (26.8%) | 318 (26.3%) | 211 (17.5%) | 44 (3.6%) |
@@
Unique topics | 219 | 216 | 220 | 221 | 226 |
Unique profiles | 1127 | 1132 | 1142 | 1143 | 1154 |
@@ 4.5 Re-identifying Users Across Sites
Finally, to respond to (Q3), we run the following re-identification experiment [6, 8]. We consider that each user visits two different websites on which two different advertisers that are colluding are embedded. For each week that we simulate, these advertisers are observing the topics of each user from their respective vantage points. Their objective is to correlate a user seen by one advertiser on the first website with the correct user seen by the other advertiser on the second website. For each user, this matching is done by returning the user(s) that have the lowest Hamming distance between the topics observed by the second advertiser and the set of topics observed by the first advertiser [8]. We leave to future work the exploration of other matching strategies. This re-identification is performed every week with only the topics that have been observed until that week for each user. We present the results of this experiment in terms of level of -anonymity that is provided to users by the API; i.e.,, means that advertisers have been able to correctly re-identify the user across their visits for that specific week, where is the number of users means that the user’s visits could not be uniquely correlated to each other, but with the visits of other users, and means that the re-identification failed. Note that the experiment is also valid, and our results exactly the same, if we were to consider instead of two advertisers only one embedded on both websites and trying to correlate visits from both vantage points. Figure 3 shows for a simulation run the cumulative distribution function of the proportion of users that are satisfied -anonymity (log scale) across the simulated weeks (recall that a topic observation for a user requires knowing their top interests for the last weeks, thus we start at week 3). Additionally, we report in Table IV the number of users that are correctly re-identified across observations for 10 simulation runs. We observe that the probability of being re-identified across visits for these users is of , , and after , , and observations, respectively. For the weeks simulated, we observe that about of the users are mapped to a group of size , which means that their cross-site visits can be correlated to each other with a higher probability than just a random guess. We conclude that some users from this real dataset are re-identified across websites through only the observations of their topics of interest in our experiment. Thus, the real users from our dataset can be fingerprinted through the Topics API. Moreover, as can be seen, the information leakage and so, privacy violation worsen over time as more users are re-identified.
Simulation | 1 observation | 2 observations | 3 observations |
Run 1 | 29 users | 28 users | 42 users |
Run 2 | 28 users | 37 users | 49 users |
Run 3 | 29 users | 47 users | 63 users |
Run 4 | 24 users | 33 users | 46 users |
Run 5 | 30 users | 39 users | 52 users |
Run 6 | 24 users | 33 users | 37 users |
Run 7 | 31 users | 38 users | 49 users |
Run 8 | 32 users | 41 users | 56 users |
Run 9 | 29 users | 27 users | 48 users |
Run 10 | 27 users | 44 users | 62 users |
mean std | 28.3 2.7 | 36.7 6.5 | 50.4 8.2 |
@@ 5 Discussion
Future Work. Overall, rigorous and thorough analyses of these new web proposals (The Privacy Sandbox for the web now encompasses more than 20 different mechanisms [33]) are needed; several interesting questions remain unsolved, such as if these proposals can deliver on their guarantees to all users, if their specification and implementation are correct, if they can be manipulated or abused, if their results can be correlated with other fingerprinting signals, if developers and users do understand how these new APIs work, etc. Regarding the Topics API, evaluating it without having access to Google’s private dataset of browsing histories can be quite challenging; we refer readers to previous analyses of the Topics API to see how prior authors (partially) got around that obstacle [5, 6]. In this paper, we call on web actors to release anonymized or synthetic datasets to help with that point. In the meantime, we envision to perform a literature review and a taxonomy of web users measurement studies, pair these results with the few datasets of browsing histories publicly available, and generate synthetic distributions of users’ topics of interest.
@@ 6 Conclusion
This paper complements previous analyses of the Topics API for the web by confirming on real data that the proposal does not provide the same privacy guarantees to all users. Specifically, we show how these real users exhibit stable and unique topics profiles and how they risk to be fingerprinted by their topics as the information leakage worsens over the time period that can be simulated with this dataset; the probability for these users to be re-identified across websites is of , , and after , , and observations of their topics, respectively. We also highlight the need for reproducible analyses of the claims made by new web proposals before their deployment. This is important to avoid replicating the same mistakes already made in the past with the same technologies being deprecated today; the introduction of e.g., web cookies later extensively abused for tracking purposes. As such, we call on web actors and industry to release representative and anonymized datasets to enable reproducible evaluations of the Topics API and other proposals from The Privacy Sandbox over time.
@@ Acknowledgments
We sincerely thank the reviewers for their constructive comments and suggestions on this paper. We also thank Josh Karlin, lead engineer from Google on the Topics API for the help provided when pointing out a subtle bug in our initial simulation for this paper; the code and quantitative results have been revised in this latest version.
@@ Appendix A Notations
Commit on Topics proposal | Commit sha of the Topics proposal studied in this paper | 2df537b (July 8, 2024) |
@@
CrUX version | Version of the CrUX top-list used in this paper | 202406 (June 2024) |
Tranco version | Version of the Tranco top-list used in this paper | G6KQK (August 3, 2024) |